# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v2

@@ 5.3. Construct Validity
5.4. Conclusion Validity
Threats to conclusion validity relate to factors, such as data inaccuracies, that can impede drawing accurate conclusions. To mitigate these threats, we adhered to best practices, including the search protocol, pilot search, and pilot data selection, as recommended by Kitchenham et al. (Keele et al., 2007) and Petersen et al. (Petersen et al., 2008b). Moreover, to further ensure the validity of our conclusions, the authors engaged in multiple brainstorming sessions to collaboratively interpret the results and finalize the conclusions.
6. Related Work
@@
6.1. Challenges, Solutions, and Performance of Multi-cloud Containers
@@
In recent years, several literature surveys have been published to analyze state-of-the-art on the applications (Pahl et al., 2017), tools and technologies (Casalicchio and Iannucci, 2020), potential and limitations (Bentaleb et al., 2022), as well as emerging trends (Watada et al., 2019) of container-based solutions for cloud computing systems, discussed below.
6.1.1. Challenges, Applications, Tools, and Emerging Trends
@@
Containers in multi-cloud environment provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms (Alonso et al., 2023c) (Saxena et al., 2021). One of the earliest SLR-based studies by Pahl et al. (Pahl et al., 2017) on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems. The SLR classified and compared the selected research studies using a conceptual framework to highlight existing trends and needs for future research. SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds. The SLR in (Alonso et al., 2023c) reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems. Similar survey studies such as (Saxena et al., 2021) concepts, challenges, requirements and future directions for multi-cloud environment are discussed. A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same (Imran et al., 2020).
6.1.2. Performance and Resource Utilization
@@
In a study by Bentaleb et al. (Bentaleb et al., 2022), the authors follow the taxonomical classification from (Pahl et al., 2017) to categorize container-based technologies for cloud systems. The study argues for the need of performance metrics to objectively define and evaluate quality attributes such as resource virtualization, service elasticity, orchestration, and multi-tenancy in containerized cloud systems. The study also highlights the needs for future research in terms of best practices, i.e., patterns and tactics (enabling reuse) and tools (supporting automation) for container-based cloud deployment. In the context of performance, the studies (Casalicchio and Iannucci, 2020) (Watada et al., 2019) report literature review and experimental analysis of factors that influence performance of container-based cloud systems. Specifically, Casalicchio et al. (Casalicchio and Iannucci, 2020) reviewed a total of 97 research studies investigating performance evaluation and run-time adaptation of container-based solution. The study highlighted several unsolved challenges such as I/O throughput optimization, performance prediction, and multi-layer monitoring performance bottlenecks. Moreover, Watada et al. (Watada et al., 2019) conducted an experimental study to compare the performance of VMs, containers and uni kernels in terms of and technological maturity using standard benchmarks and observed containers to optimise performance of container-based. The performance of containerized clouds is also determined by resource utilization, which includes but is not limited to CPU utilization, memory footprints, energy consumption, network bandwidth, and execution time.
@@
Kapil et al. (Netaji and Bhole, 2021) reviewed 64 research papers to gain insight into resource allocation, management, and scheduling. Furthermore, limitations of existing resource allocation algorithms are discussed, indicating the need to investigate algorithms or techniques of performance optimization of containers for cloud systems. The study by Maenhaut et al. (Maenhaut et al., 2020) provided an overview of the current state of the art regarding resource management within the broad sense of cloud computing, complementary to existing surveys in the literature. The study investigated how research is adapting to the recent evolution within cloud solutions, including container technologies.
6.2. Container-based Cloud Orchestration, Deployment, and Security
@@ 6.2.1. Orchestration and Deployment
Naweiluo et al. (Zhou et al., 2022) explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing. The results of this study indicate that containers enhance application deployment efficiency, but face challenges in HPC due to high security levels and the need for extensive libraries and packages that affect cloud portability, often resulting in vendor lock-ins. The study also provides a survey and taxonomy on containerization and orchestration efforts in HPC, pointing out differences with cloud environments and identifying future research and engineering potentials. A systematic mapping study by Naylor et al.(Bachiega et al., 2018) explores virtualization in container-based cloud deployments. The mapping study, based on a comprehensive review of major databases, identifies a significant research gap in the performance evaluation of containers, underscoring the need for further investigation on cloud deployment. Carlos et al. (Awada, 2018) present an optimization strategy for deploying microservices in multi-cloud environment focusing on minimizing cloud service costs, network latency, and startup time for new microservices. Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II. This highlights its effectiveness in container and VM orchestration, offering a significant enhancement in deployment of multi-cloud solution in containers. Emiliano et al. (Paladi et al., 2018) survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field. Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively. Uchechukwu Awada (Guerrero et al., 2018b) reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos. The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.
@@
Koustabh et al. (Dolui and Kiraly, 2018) explore multi-container deployment on IoT gateways to meet the stringent latency requirements of advanced IoT applications. Through their study within the AGILE project, they highlight containerization’s role in overcoming the diversity and resource constraints of IoT gateways, showcasing containerization’s advantages for IoT gateway performance optimization. Their research underscores the potential of containerized environments in enhancing application compatibility, portability, and efficient deployment across diverse hardware architectures. Matteo et al. (Nardelli et al., 2018) introduce the Adaptive Container Deployment (ACD) model to optimize containerized application deployment in geo-distributed environments, focusing on IoT and fog computing resources. ACD, formulated as an Integer Linear Programming problem, aims to improve application performance by leveraging containers’ horizontal and vertical elasticity. The study evaluates ACD’s effectiveness against greedy heuristics, highlighting the need for advanced orchestration solutions to exploit emerging computing environments’ characteristics efficiently.
6.2.2. Security of Containerized Clouds
@@
Nicolae et al. (Casalicchio, 2019) analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework. Their research identifies potential attack scenarios and security enforcement mechanisms, aiming to enhance security guarantees for cloud operators in a multi-cloud setting. Mohammad et al. (Bokhari et al., 2018) offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models. They discuss the security requirements for public and private clouds across different service models, aiming to provide researchers with a comprehensive understanding of cloud computing’s potential and security challenges. Sari et al. (Sultan et al., 2019) survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape. Their analysis covers both software-based solutions utilizing Linux kernel features and hardware-based solutions for enhanced security. The study aims to clarify container security requirements and encourage further research in addressing potential vulnerabilities and attacks. Hendrik et al. (Graupner et al., 2015b) tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective. They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control. Rohan et al. (Gupta et al., 2016) propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption. This approach aims to secure data while optimizing resource usage and cost, highlighting the benefits of container technology in improving cloud computing security and efficiency.
6.3. Conclusive Summary
@@
Current survey-based studies and systematic reviews, discussed above (e.g., (Casalicchio and Iannucci, 2020) (Pahl et al., 2017)), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems. These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends. Table 16 presents an overview of the differences between this study and existing secondary studies. The symbols used in the table are as follows:  indicates “Yes”,  indicates “No”, and $\triangle$ indicates “Partially”. After thoroughly comparing our results with existing studies (e.g., (Pahl et al., 2017), (Casalicchio and Iannucci, 2020), (Netaji and Bhole, 2021)), it is clear that none of the results are fully identical with our findings. Our study distinguishes itself from prior research in several ways. First, our work systematically categorizes container utilization in multi-cloud environment into coherent categories and themes. Some of the key themes are Scalability and High Availability, Performance and Optimization, Security and Privacy, and Multi-Cloud Container Monitoring and Adaptation. We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided. Furthermore, we explored the QAs and associated tactics for containerization in multi-cloud environment. This contribution offers practitioners actionable insights into tailoring their containerization strategies to meet desired quality standards. Lastly, our study introduces four distinct frameworks consisting of challenges and solutions related to security, automation, deployment, and monitoring within the context of containerization in multi-cloud environment.
@@
Table 16. Comparison of This SLR Results with Existing Secondary Studies. MS indicate “Mix Study”, LR indicate “Literature Review”  indicates “Yes”,  indicates “No,” and $\triangle$ indicates “Partially.”
Contributions This Study (Pahl et al., 2017) (Casalicchio and Iannucci, 2020) (Netaji and Bhole, 2021) (Bentaleb et al., 2022) (Watada et al., 2019) (Saxena et al., 2021) (Imran et al., 2020) (Casalicchio, 2019)
@@
7. Conclusions
This SMS presents the current state of research regarding containerization in multi-cloud environment, focusing on aims of selected studies, the roles of containers, containers implementation strategies, architectural patterns and strategies, quality attributes and corresponding tactics. Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment. From an initial pool of extensive search results, 86 relevant studies were finally chosen to answer the research questions presented in Table 1. The key findings of this SMS are listed below.
@@ 7. Conclusions
  • This SMS presents a comprehensive security challenge-solution framework for containerized applications in multi-cloud environment. This framework offers valuable insights and actionable strategies for practitioners and researchers to enhance the security of containerized applications, address complex multi-cloud security concerns, and ensure compliance with various regulatory requirements, thereby enabling the development of more robust and secure containerized applications in multi-cloud environment.
  • This SMS presents a catalogs for the automation challenge-solution catalogs for containerized applications in multi-cloud environment. The automation challenge-solution catalog offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environment. By offering specific solutions for each challenge, this catalog aids in the development of more efficient, secure, and adaptable automation processes, empowering practitioners to streamline their operations and enabling researchers to advance the field of multi-cloud containerization.
@@
The dataset of this study has been provided online: https://zenodo.org/record/10732611.
@@ Acknowledgments
This research is funded by Business Finland through QLEAP (2022-24) project, and by the NSFC China under Grant No. 62172311.
8. Appendices
{xltabular}
—c—p4.5cm—p4.5cm—c—c— List of selected studies for containerization in multi-cloud environments
Study ID Title Venue Year Reference

### v3

@@ 5.3 Construct Validity
5.4 Conclusion Validity
Threats to conclusion validity relate to factors, such as data inaccuracies, that can impede drawing accurate conclusions. To mitigate these threats, we adhered to best practices, including the search protocol, pilot search, and pilot data selection, as recommended by Kitchenham et al. [17] and Petersen et al. [180]. Moreover, to further ensure the validity of our conclusions, the authors engaged in multiple brainstorming sessions to collaboratively interpret the results and finalize the conclusions.
6 Related Work
@@
6.1 Challenges, Solutions, and Performance of Multi-cloud Containers
@@
In recent years, several literature surveys have been published to analyze state-of-the-art on the applications [183], tools and technologies [184], potential and limitations [185], as well as emerging trends [186] of container-based solutions for cloud computing systems, discussed below.
6.1.1 Challenges, Applications, Tools, and Emerging Trends
@@
Containers in multi-cloud environment provide a standardized and portable way to package, deploy, and run applications, ensuring seamless deployment and management across diverse cloud platforms [187] [188]. One of the earliest SLR-based studies by Pahl et al. [183] on containerized clouds reviewed a total of 46 studies to identify, taxonomically classify and systematically compare the existing research on containers and their application in the context of cloud-based systems. The SLR classified and compared the selected research studies using a conceptual framework to highlight existing trends and needs for future research. SLR results indicate container orchestration, microservice delivery, continuous development and deployment as emerging trends of research on containerized clouds. The SLR in [187] reviewed a total of 88 studies (published from 2011 - 2021) on multi-cloud systems. Similar survey studies such as [188] concepts, challenges, requirements and future directions for multi-cloud environment are discussed. A survey of existing approaches and solutions provided by different multi-cloud architectures is entailed along with analysis of the pros and cons of different architectures while comparing the same [189].
6.1.2 Performance and Resource Utilization
@@
In a study by Bentaleb et al. [185], the authors follow the taxonomical classification from [183] to categorize container-based technologies for cloud systems. The study argues for the need of performance metrics to objectively define and evaluate quality attributes such as resource virtualization, service elasticity, orchestration, and multi-tenancy in containerized cloud systems. The study also highlights the needs for future research in terms of best practices, i.e., patterns and tactics (enabling reuse) and tools (supporting automation) for container-based cloud deployment. In the context of performance, the studies [184] [186] report literature review and experimental analysis of factors that influence performance of container-based cloud systems. Specifically, Casalicchio et al. [184] reviewed a total of 97 research studies investigating performance evaluation and run-time adaptation of container-based solution. The study highlighted several unsolved challenges such as I/O throughput optimization, performance prediction, and multi-layer monitoring performance bottlenecks. Moreover, Watada et al. [186] conducted an experimental study to compare the performance of VMs, containers and uni kernels in terms of and technological maturity using standard benchmarks and observed containers to optimise performance of container-based. The performance of containerized clouds is also determined by resource utilization, which includes but is not limited to CPU utilization, memory footprints, energy consumption, network bandwidth, and execution time.
@@
Kapil et al. [190] reviewed 64 research papers to gain insight into resource allocation, management, and scheduling. Furthermore, limitations of existing resource allocation algorithms are discussed, indicating the need to investigate algorithms or techniques of performance optimization of containers for cloud systems. The study by Maenhaut et al. [191] provided an overview of the current state of the art regarding resource management within the broad sense of cloud computing, complementary to existing surveys in the literature. The study investigated how research is adapting to the recent evolution within cloud solutions, including container technologies.
6.2 Container-based Cloud Orchestration, Deployment, and Security
@@ 6.2.1 Orchestration and Deployment
Naweiluo et al. [192] explored containerization in High Performance Computing (HPC) environments, contrasting it with cloud computing. The results of this study indicate that containers enhance application deployment efficiency, but face challenges in HPC due to high security levels and the need for extensive libraries and packages that affect cloud portability, often resulting in vendor lock-ins. The study also provides a survey and taxonomy on containerization and orchestration efforts in HPC, pointing out differences with cloud environments and identifying future research and engineering potentials. A systematic mapping study by Naylor et al.[193] explores virtualization in container-based cloud deployments. The mapping study, based on a comprehensive review of major databases, identifies a significant research gap in the performance evaluation of containers, underscoring the need for further investigation on cloud deployment. Carlos et al. [194] present an optimization strategy for deploying microservices in multi-cloud environment focusing on minimizing cloud service costs, network latency, and startup time for new microservices. Their approach uses a Non-dominated Sorting Genetic Algorithm II (NSGA-II) compared to a Greedy First-Fit algorithm, demonstrating a 300% improvement with NSGA-II. This highlights its effectiveness in container and VM orchestration, offering a significant enhancement in deployment of multi-cloud solution in containers. Emiliano et al. [195] survey container orchestration, proposing a reference architecture for autonomic orchestrators, and identifying research challenges in the field. Their work emphasizes the importance of container technologies in cloud environments and the need for advanced orchestration solutions to manage complex multi-container applications effectively. Uchechukwu Awada [196] reviews container orchestration tools and platforms, comparing the architectures, components, and capabilities of several Container Service Platforms (CSPs) and orchestration tools like Amazon ECS, Kubernetes, Docker Swarm, and Mesos. The study offers insights into the current state of container orchestration and suggests future research directions, serving as a guide for developers and organizations.
@@
Koustabh et al. [197] explore multi-container deployment on IoT gateways to meet the stringent latency requirements of advanced IoT applications. Through their study within the AGILE project, they highlight containerization’s role in overcoming the diversity and resource constraints of IoT gateways, showcasing containerization’s advantages for IoT gateway performance optimization. Their research underscores the potential of containerized environments in enhancing application compatibility, portability, and efficient deployment across diverse hardware architectures. Matteo et al. [198] introduce the Adaptive Container Deployment (ACD) model to optimize containerized application deployment in geo-distributed environments, focusing on IoT and fog computing resources. ACD, formulated as an Integer Linear Programming problem, aims to improve application performance by leveraging containers’ horizontal and vertical elasticity. The study evaluates ACD’s effectiveness against greedy heuristics, highlighting the need for advanced orchestration solutions to exploit emerging computing environments’ characteristics efficiently.
6.2.2 Security of Containerized Clouds
@@
Nicolae et al. [199] analyze security challenges in cloud orchestration for multi-cloud deployments, proposing a security-enabled orchestration framework. Their research identifies potential attack scenarios and security enforcement mechanisms, aiming to enhance security guarantees for cloud operators in a multi-cloud setting. Mohammad et al. [200] offer a detailed analysis of cloud computing, defining its essential characteristics, architecture, service models (SaaS, PaaS, IaaS), and deployment models. They discuss the security requirements for public and private clouds across different service models, aiming to provide researchers with a comprehensive understanding of cloud computing’s potential and security challenges. Sari et al. [201] survey the landscape of container security, addressing challenges and solutions across four generalized use cases within the host-container threat landscape. Their analysis covers both software-based solutions utilizing Linux kernel features and hardware-based solutions for enhanced security. The study aims to clarify container security requirements and encourage further research in addressing potential vulnerabilities and attacks. Hendrik et al. [202] tackle privacy, security, and trust issues in cloud computing through a multi-cloud architecture perspective. They propose a novel technique for enhancing security and unifying access control mechanisms across cloud providers, addressing the challenges of inadequate cross-provider APIs and non-unified access control. Rohan et al. [203] propose a security-enhancing methodology for cloud data storage using container clustering and Docker instances for on-demand encryption. This approach aims to secure data while optimizing resource usage and cost, highlighting the benefits of container technology in improving cloud computing security and efficiency.
6.3 Conclusive Summary
@@
Current survey-based studies and systematic reviews, discussed above (e.g., [184] [183]), aim to consolidate the latest findings in published research concerning container-based solutions for cloud computing systems. These studies investigated various aspects of multi-cloud, including exploring applications, tools, and technologies, assessing potential and limitations, and identifying emerging trends. Table 16 presents an overview of the differences between this study and existing secondary studies. The symbols used in the table are as follows: ✗ indicates “Yes”, ✓ indicates “No”, and $\triangle$ indicates “Partially”. This study distinguishes itself from prior secondary studies in multiple critical ways, thereby advancing the state of the art in the domain of containerization for multi-cloud environments. While existing surveys and mapping studies (e.g., [183], [184], [190]) provide valuable overviews of tools, technologies, performance, and resource management, they often focus on isolated aspects or lack integration across key quality and deployment dimensions. In contrast, our SMS presents a comprehensive, theme-based classification of 121 selected studies—making it the most extensive secondary analysis in this domain to date. First, our work systematically categorizes container utilization in multi-cloud environment into coherent categories and themes. Some of the key themes are Scalability and High Availability, Performance and Optimization, Security and Privacy, and Multi-Cloud Container Monitoring and Adaptation. We identified and classified seventy-four patterns and strategies for containerization, which existing studies have not provided. This level of operational detail is absent in prior work. Furthermore, we explored the QAs and associated tactics for containerization in multi-cloud environment. Unlike existing reviews that superficially mention quality attributes, we provide a dedicated framework linking QAs with implementation tactics, offering practical insights for engineering design decisions—an aspect previously unexplored. Our SMS also identifies and formalizes four distinct challenge-solution frameworks in the areas of security, automation, deployment, and monitoring. These frameworks synthesize fragmented research findings into cohesive, actionable knowledge structures, not previously consolidated in the literature. Lastly, our comparative analysis (Table 16) demonstrates that none of the previous studies collectively cover the breadth and depth addressed by our SMS, particularly in integrating performance, security, orchestration, and QA strategies across the entire lifecycle of container-based multi-cloud systems. By addressing these multiple dimensions holistically, our study not only fills significant gaps in the literature but also provides a practical roadmap for researchers and practitioners aiming to optimize container deployments in complex, heterogeneous cloud environments.
@@
Table 16: Comparison of this SLR results with existing secondary studies. MS indicates “Mix Study”, LR indicates “Literature Review”, ✓ indicates “Yes”, ✗ indicates “No”, and $\triangle$ indicates “Partially”
Contributions This Study [183] [184] [190] [185] [186] [188] [189] [199]
@@
7 Conclusions
This SMS presents the current state of research regarding containerization in multi-cloud environment, focusing on aims of selected studies, the roles of containers, containers implementation strategies, architectural patterns and strategies, quality attributes and corresponding tactics. Additionally, this SMS also explores deployment, monitoring, security challenges, accompanied by their respective solutions, tools, and frameworks used to implement containerized applications in multi-cloud environment. We investigated 121 relevant studies to answer the RQs in Table 1 with key findings of this SMS include:
@@ 7 Conclusions
  • This SMS presents a comprehensive security challenge-solution framework for containerized applications in multi-cloud environment. This framework offers valuable insights and actionable strategies for practitioners and researchers to enhance the security of containerized applications, address complex multi-cloud security concerns, and ensure compliance with various regulatory requirements, thereby enabling the development of robust and secure containerized applications for multi-clouds.
  • This SMS presents a catalogs for the automation challenge-solution catalogs for containerized applications in multi-cloud environment. The automation challenge-solution catalog offers a comprehensive guide for practitioners and researchers to address the complex challenges associated with automating various aspects of deploying, managing, scaling, testing, migrating, and ensuring interoperability of containerized applications in multi-cloud environments.
@@
Link to our dataset is in the reference [11].
@@ Acknowledgments
This research is funded by Business Finland through QLEAP (2022-24) project, the National Natural Science Foundation of China (NSFC) under Grant No. 62172311, and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.
Declaration of AI Assistance